# Distribution-aware Fairness Learning in Medical Image Segmentation From A Control-Theoretic Perspective

### v1

@@ 4.4 Results
We first evaluate dMoE for 2D segmentation on benchmark datasets. For the Harvard-FairSeg neuroretinal rim and optic nerve segmentation tasks, our method consistently achieves SOTA performance in terms of ES-Dice and ES-IoU, surpassing previous fairness learning approaches, as shown in Table 1. Specifically, our method demonstrates superior performance for the minor attribute subgroup (Black), achieving a Dice score of 0.776 for rim segmentation, compared to FairDiff (0.743) and FEBS (0.733). Similarly, for optic cup segmentation, our method significantly improves performance for the Asian attribute subgroup, achieving a Dice score of 0.844 compared to FairDiff (0.832) and FEBS (0.825). Although MoE, which removes the distribution-aware gating network from our proposed method, achieves comparable performance (0.845), dMoE outperforms MoE across other subgroups for both rim and cup segmentation. Furthermore, the shared structure within dMoe module across subgroups enables synergistic effects, maintaining or even improving performance for the major attribute subgroup (White). For the HAM10000 skin lesion segmentation task, our method also demonstrates SOTA performance in ES-Dice and ES-IoU, with scores of 0.841 and 0.749, respectively, outperforming MoE (0.836 and 0.742) and FEBS (0.787 and 0.679), as shown in Table 2. Specifically, loss function modification methods, such as FEBS, exhibit inferior performance across attribute subgroups compared to the baseline method. This suggests that when the training dataset for each subgroup is insufficient, these methods may sacrifice overall performance. MoE shows promising results for minor attribute subgroups, dMoE further enhances equity across all subgroups.
@@
We further evaluate dMoE for 3D radiotherapy target segmentation, using T-stage as a distributional attribute. Table 3 summarizes the segmentation performance. Despite the test set being acquired from a different hospital with a distinct data distribution, dMoE demonstrates promising generalization performance. The strength of dMoE is particularly evident in underrepresented subgroups, such as T1 and T4 where the frequency is significantly lower than that of T2 and T3, by resulting the most notable performance improvements in these subgroups. Specifically, dMoE demonstrates robust and consistent performance, particularly excelling in the T4 subgroup with a Dice score of 0.767, significantly outperforming FEBS (0.660) and MoE (0.683). Moreover, despite the limited sample size of the T1 subgroup, dMoE achieves superior performance. This consistent improvement demonstrates dMoE’s generalizability and robustness, even when applied to datasets from different centers with varying data distributions.
@@ 4.4.3 Ablation Studies
We further conduct additional analyses to explore the effects of various configurations of the dMoE module. Specifically, we vary the placement of the dMoE module within the encoder and decoder and investigate whether its parameters should be shared or separated across layers to determine the optimal architecture. Detailed ablation study results are provided in Section A.4.
@@ Appendix A Appendix
Testset 2000 169 299 1532 1061 121 469 328 120 24 132 1 55 64 12
A.3 Comparison on Computations.
We compare the computational complexity of incorporating the MoE or dMoE module into the backbone in Table 7.
Table 7: Computational complexity comparison.
@@
A.4 Ablation Study Results.
@@
We perform ablation studies to examine the effects of various configurations of the dMoE module. Specifically, we investigate performance changes based on 1) dMoE Location within the network, by inserting the dMoE module at different layers within the encoder, the decoder, or both for comparison. We also examine whether 2) dMoE Parameters should be shared or separated across layers to determine the optimal architecture. In the case of parameter sharing for CNNs, we incorporate additional linear layers before and after the dMoE module to match the channel dimensions across different layer blocks. For performing experiments, we use the HAM10000 dataset for Transformer-based architecture (TransUNet), while for CNN-based architecture (3D ResUNet), we utilize a radiotherapy target segmentation task, by reporting performance using the ES-Dice and Dice score as the metric.
@@
As shown in Table 8 and Table 9, inserting the dMoE module within the encoder, decoder, or both yields comparable performance for both architectures, particularly for minority groups (The subgroup with Age $<$ 20 for HAM10000 dataset and the T4-stage subgroup for radiotherapy target dataset). Therefore, we retain the dMoE module within the encoder only to optimize network training efficiency. In the Transformer-based architecture, sharing dMoE parameters across layers enhances performance due to consistent dimensionality between layers. Conversely, in CNN-based architecture, differences in layer-wise dimensionality reduce the effectiveness of parameter sharing. This necessitates adding dimension-matching linear layers before and after the dMoE modules; however, performance remains comparable. As a result, we adopt a layer-wise dMoE configuration specifically tailored for CNNs.

### v2

@@ 4.4 Results
We first evaluate dMoE for 2D segmentation on benchmark datasets. For the Harvard-FairSeg neuroretinal rim and optic nerve segmentation tasks, our method consistently achieves SOTA performance in terms of ES-Dice and ES-IoU, surpassing previous fairness learning approaches, as shown in Table 1. Specifically, our method demonstrates superior performance for the minor attribute subgroup (Black), achieving a Dice score of 0.776 for rim segmentation, compared to FairDiff (0.743) and FEBS (0.733). Similarly, for optic cup segmentation, our method significantly improves performance for the Asian attribute subgroup, achieving a Dice score of 0.844 compared to FairDiff (0.832) and FEBS (0.825). Although MoE, which removes the distribution-aware gating network from our proposed method, achieves comparable performance (0.845), dMoE outperforms MoE across other subgroups for both rim and cup segmentation. Furthermore, the shared structure within dMoe module across subgroups enables synergistic effects, maintaining or even improving performance for the major attribute subgroup (White). For the HAM10000 skin lesion segmentation task, our method also demonstrates SOTA performance in ES-Dice and ES-IoU, with scores of 0.801 and 0.725, respectively, outperforming MoE (0.796 and 0.721) and FEBS (0.757 and 0.667), as shown in Table 2. Specifically, loss function modification methods, such as FEBS, exhibit inferior performance across attribute subgroups compared to the baseline method. This suggests that when the training dataset for each subgroup is insufficient, these methods may sacrifice overall performance. While MoE demonstrates promising performance for minor attribute subgroups, dMoE further improves fairness.
@@
We further evaluate dMoE for 3D radiotherapy target segmentation, using T-stage as a distributional attribute. Table 3 summarizes the segmentation performance. Despite the test set being acquired from a different hospital with a distinct data distribution, dMoE demonstrates promising generalization performance. The strength of dMoE is particularly evident in underrepresented subgroups, such as T1 and T4 where the frequency is significantly lower than that of T2 and T3, by resulting the most notable performance improvements in these subgroups. Specifically, dMoE demonstrates robust and consistent performance, particularly excelling in the T4 subgroup with a Dice score of 0.778, significantly outperforming FEBS (0.685) and MoE (0.708). Moreover, despite the limited sample size of the T1 subgroup, dMoE achieves superior performance. This consistent improvement demonstrates dMoE’s generalizability and robustness, even when applied to datasets from different centers with varying data distributions.
@@
We also present a detailed comparison between the dMoE attribute-wise gating mechanism and multiple independently trained networks in terms of computational complexity and performance, as shown in Table 8 of Section A.3. The results demonstrate that dMoE not only achieves higher segmentation accuracy but also maintains greater computational efficiency, outperforming specialized models with a single unified network.
4.4.3 Radiotherapy Target Segmentation with Diverse Clinical Attributes
In clinical practice, radiotherapy targets are determined not solely by tumor stage or visible tumor imaging but through the integration of anatomical imaging with clinical parameters such as T-stage, Gleason Grade Group (GG), and prostate-specific antigen (PSA) levels to account for potential microscopic spread (National Comprehensive Cancer Network, 2024). To evaluate the broader applicability of our method, we further tested additional clinical factors which reflect pathological tumor differentiation but do not indicate the precise tumor location or extent. As shown in Section A.4, our results demonstrate fairness improvements comparable to those observed with T-stage. These finding show that incorporating clinical indicators—even when not directly related to visible anatomical structures—can enhance segmentation accuracy across subpopulations.
@@ 4.4.4 Ablation Studies
We further conduct additional analyses to explore the effects of various configurations of the dMoE module in Section A.5. Specifically, we vary the placement of the dMoE module within the encoder and decoder and investigate whether its parameters should be shared or separated across layers to determine the optimal architecture. We further ablate each component of the optimal control framework to assess the contribution of control theory to distribution-aware fairness learning. Detailed results of the ablation study are provided in Table 11 and Table 12.
@@ 5 Conclusion, Limitations & Future Work
Acknowledgements
This work was supported by the National Institutes of Health, USA R01HL159183, and also supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2024-00345854).
@@ Appendix A Appendix
Testset 2000 169 299 1532 1061 121 469 328 120 24 275 11 129 114 21
A.3 Comparison on Computational Efficiency.
We compare the computational complexity of incorporating the dMoE module into the backbone in Table 8. Additionally, Table 8 presents a comparison between the dMoE attribute-wise gating mechanism and multiple independently trained networks, in terms of both computational complexity and performance, focusing on the 3D radiotherapy target segmentation task using tumor stage as the distribution attribute. The results show that dMoE achieves superior performance while maintaining greater computational efficiency, outperforming individually specialized networks with a single unified model.
@@
Table 7: Computational complexity comparison.
Method GFlops $\downarrow$ All T1 T2 T3 T4
(n=275) (n=11) (n=129) (n=114) (n=21)
ES-Dice(D) Dice Dice Dice Dice Dice
dMoE (Ours) 1761.30 0.499 0.650 0.718 0.585 0.693 0.778
Multiple networks for each attribute 5729.44 0.457 0.606 0.599 0.515 0.681 0.760
Table 8: Comparison to multiple networks for each attribute.
A.4 Further Clinical Attribute Analysis on Radiotherapy Target Segmentation.
To further evaluate the ability of the proposed dMoE module to capture the diversity of real-world clinical settings, we incorporate an additional clinical parameter as a distribution attribute in the prostate cancer study. Specifically, we use the Gleason Grade Group (GG) and prostate-specific antigen (PSA) level to represent pathological differentiation and disease progression, respectively. As shown in Table 9, our method demonstrates robust performance across various subgroups, particularly in underrepresented groups such as GG 6, 9, and 10. Similarly, Table 10 shows that the proposed dMoE module enhances equity across PSA subgroups, with notable improvements in underrepresented groups such as PSA levels 0 and 4.
Table 9: Comparison on 3D radiotherapy target segmentation with Gleason Grade Groups (GG) as the distribution attribute.
Method All (n=275) GG level 6 (n=31) GG 7 (n=125) GG 8 (n=62) GG 9 (n=47) GG 10 (n=10)
ES-Dice Dice ES-IoU IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU
3D ResUNet (Çiçek et al., 2016) 0.512 0.610 0.389 0.461 0.562 0.417 0.578 0.429 0.650 0.501 0.669 0.517 0.623 0.474
+ FEBS (Tian et al., 2024) 0.451 0.593 0.337 0.441 0.501 0.349 0.557 0.406 0.628 0.473 0.686 0.534 0.650 0.494
+ MoE 0.447 0.608 0.341 0.461 0.514 0.366 0.565 0.419 0.653 0.505 0.704 0.559 0.689 0.536
+ dMoE 0.473 0.638 0.361 0.494 0.672 0.533 0.566 0.419 0.657 0.511 0.750 0.614 0.750 0.612
Table 10: Comparison on 3D radiotherapy target segmentation with prostate-specific antigen (PSA) level as the distribution attribute.
Method All (n=275) PSA level 0 (n=51) level 1 (n=84) level 2 (n=60) level 3 (n=29) level 4 (n=52)
ES-Dice Dice ES-IoU IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU
3D ResUNet (Çiçek et al., 2016) 0.504 0.608 0.380 0.460 0.552 0.402 0.578 0.429 0.643 0.498 0.643 0.495 0.659 0.508
+ FEBS (Tian et al., 2024) 0.561 0.588 0.412 0.434 0.571 0.422 0.584 0.429 0.609 0.455 0.592 0.441 0.585 0.427
+ MoE 0.461 0.606 0.348 0.459 0.532 0.383 0.553 0.407 0.632 0.486 0.667 0.519 0.709 0.563
+ dMoE 0.575 0.654 0.442 0.510 0.664 0.530 0.612 0.465 0.643 0.494 0.660 0.513 0.723 0.582
A.5 Ablation Study Results.
@@
We perform ablation studies to examine the effects of various configurations of the dMoE module. Specifically, we investigate performance changes based on 1) dMoE Location within the network, by inserting the dMoE module at different layers within the encoder, the decoder, or both for comparison. We also examine whether 2) dMoE Parameters should be shared or separated across layers to determine the optimal architecture. In the case of parameter sharing for CNNs, we incorporate additional linear layers before and after the dMoE module to match the channel dimensions across different layer blocks. We further analyze the contribution of 3) Optimal Control Components by defining the mode-switching feedback control as the dMoE module. To assess its impact, we ablate both the mode-switching and feedback control components. For performing experiments, we use the HAM10000 dataset for Transformer-based architecture (TransUNet), while for CNN-based architecture (3D ResUNet), we utilize a radiotherapy target segmentation task, by reporting performance using the ES-Dice and Dice score as the metric.
@@
As shown in Table 11 and Table 12, inserting the dMoE module within the encoder, decoder, or both yields comparable performance for both architectures, particularly for minority groups (The subgroup with Age $<$ 20 for HAM10000 dataset and the T4-stage subgroup for radiotherapy target dataset). Therefore, we retain the dMoE module within the encoder only to optimize network training efficiency. In the Transformer-based architecture, sharing dMoE parameters across layers enhances performance due to consistent dimensionality between layers. Conversely, in CNN-based architecture, differences in layer-wise dimensionality reduce the effectiveness of parameter sharing. This necessitates adding dimension-matching linear layers before and after the dMoE modules; however, performance remains comparable. As a result, we adopt a layer-wise dMoE configuration specifically tailored for CNNs. In the analysis of optimal control components, a naive adaptation of feedback control resulted in marginal improvements or even performance degradation depending on the dataset. In contrast, our proposed mode-switching control consistently improved both ES-Dice and Dice scores across experiments.
@@
3) Optimal Control Components Control Mode-switching
Feedback ✓(Ours) 0.841 0.884 0.864 0.881 0.890 0.901 0.880
Feedback 0.836 0.882 0.864 0.875 0.889 0.904 0.882
Non-feedback 0.826 0.882 0.863 0.871 0.897 0.906 0.882
@@
3) Optimal Control Components Control Mode-switching
Feedback ✓(Ours) 0.499 0.650 0.718 0.585 0.693 0.778
Feedback 0.451 0.608 0.492 0.542 0.674 0.708
Non-feedback 0.509 0.615 0.524 0.573 0.668 0.637